OpenStudy (anonymous):

Determine whether the surfaces x^2+8x+y^2-9y+z^2-6z+65=0 and x^2+y^2=3x-2y intersect. do i use vectors?

myininaya (myininaya):

I would try writing it in this form first (x-h)^2+(y-k)^2=r^2

myininaya (myininaya):

both of them

myininaya (myininaya):

\[x^2+8x+4^2+y^2-9y+(\frac{9}{2})^2+z^2-6z+3^2=-65+4^2+(\frac{9}{2})^2+3^2\]

myininaya (myininaya):

i didn't realize that one was a sphere until i seen all the z's oh i got cut off

OpenStudy (anonymous):

oh shoot sorry, it's not +65, it's -85 at the end

myininaya (myininaya):

oh no i didn't get cut off

OpenStudy (anonymous):

is the 2nd equation still a sphere? becasue there aren't any z components

myininaya (myininaya):

\[x^2+8x+4^2+y^2-9y+(\frac{9}{2})^2+z^2-6z+3^2=-85+4^2+(\frac{9}{2})^2+3^2\]

myininaya (myininaya):

its a cirlce

myininaya (myininaya):

is this correct?

OpenStudy (anonymous):

ohh ok. yeah that makes sense

myininaya (myininaya):

so i have =-85 not =85 right?

myininaya (myininaya):

\[(x+4)^2+(y-\frac{9}{2})^2+(z-3)^2=-85+16+\frac{81}{4}+9\]

OpenStudy (anonymous):

well it should be ....\[z^{2}\]-6x-85, so it should be =85 does that make sense?

myininaya (myininaya):

ok cool

myininaya (myininaya):

\[(x+4)^2+(y-\frac{9}{2})^2+(z-3)^2=85+16+\frac{81}{4}+9\] so that means we have this instead

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

should i be setting the radii equal to eachother? or using parametrics?

myininaya (myininaya):

\[(x+4)^2+(y-\frac{9}{2})^2+(z-3)^2=130.25\]

myininaya (myininaya):

i haven't thought that far ahead

myininaya (myininaya):

i'm still thinking as i go

myininaya (myininaya):

\[x^2-3x+y^2+2y=0\]

myininaya (myininaya):

so this is that second equation

OpenStudy (anonymous):

yeah

myininaya (myininaya):

\[x^2-3x+(\frac{3}{2})^2+y^2-2y+1^2=0+(\frac{3}{2})^2+1^2\]

myininaya (myininaya):

\[(x-\frac{3}{2})^2+(y-1)^2=\frac{9}{4}+1\]

myininaya (myininaya):

\[(x-\frac{3}{2})^2+(y-1)^2=3.25\]

myininaya (myininaya):

so what to do next....thinking....

OpenStudy (binary3i):

justdo the cros product of the vectors of the planes if it trunsout to be0 then they dont intersect

OpenStudy (anonymous):

ohhh ok thanks. that makes a little more sense.

myininaya (myininaya):

Show me how to find the vector of the plane. I want to see.

OpenStudy (anonymous):

hmmm i thought i had it, but looking i;m not finding what i want in my text book

myininaya (myininaya):

Do you know how to write the vectors?

OpenStudy (anonymous):

yeah, but not from that equation. just from points or components given.

myininaya (myininaya):

I think I was going the wrong way about it

myininaya (myininaya):

earlier

OpenStudy (anonymous):

oh, by completing the square?

myininaya (myininaya):

write now i'm looking at pauls notes and you were write we need to write the equations in parametric form

OpenStudy (anonymous):

gahhhhhhh

myininaya (myininaya):

lol right binary to find how how to write the vectors we need to write the equation in parametric form?

myininaya (myininaya):

well maybe we need the radius and putting it in that form does give us the radius

myininaya (myininaya):

so maybe it wasn't a total waste

OpenStudy (binary3i):

take two points on the circle and find their vectors from center. do this for both the circles and go for crossproduct.

OpenStudy (binary3i):

this will give the plane vector

OpenStudy (binary3i):

if dcXio gives you some thing then it means the plane intersect but for circles it shoud be worked out.